OpenStudy (anonymous):

What is the value of the discriminant in the equation shown below? x^2+3x-6=0 What is the value of the discriminant in the equation shown below? x^2+3x-6=0 @Mathematics

jimthompson5910 (jim_thompson5910):

D = b^2 - 4ac D = 3^2 - 4(1)(-6) D = 9 +24 D = 33 So the discriminant is 33

OpenStudy (anonymous):

33

jimthompson5910 (jim_thompson5910):

This tells us that the polynomial has 2 real solutions (and they are irrational because 33 is NOT a perfect square)

OpenStudy (anonymous):

yeah thats true